OpenStudy (anonymous):

Which statement is true for the equation 2x - 2x - 7 = -7? A. It has infinitely many solutions B. It has two solutions C. It has one solution D. it has no solution

pooja195 (pooja195):

Have you solved it yet?

OpenStudy (anonymous):

no

pooja195 (pooja195):

Solve it please.

pooja195 (pooja195):

2x - 2x - 7 = -7 Combine like terms 2-2=?

OpenStudy (anonymous):

zero

pooja195 (pooja195):

Good now we are trying to isolate x so we add 7 to both sides -7+7=?

pooja195 (pooja195):

?

pooja195 (pooja195):

7-7=?

OpenStudy (anonymous):

zero

pooja195 (pooja195):

Good so when we have 0=0 or a number = the same number it would be infinte solutions

OpenStudy (anonymous):

ohh okayy thank youu

pooja195 (pooja195):

:)